OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. sin(pi/2)cos(pi/7)+cos(pi/2)sin(pi/7)

OpenStudy (anonymous):

@Hero

hero (hero):

Use angle addition formulas

OpenStudy (anonymous):

angle addition formulas?

OpenStudy (anonymous):

@Hero which one?

OpenStudy (anonymous):

so it'd be sin((pi/2)(pi/2))? I don't get it

zepdrix (zepdrix):

So we want to apply the Sine Angle Addition formula: \[\Large\rm \sin(\color{royalblue}{a})\cos(\color{orangered}{b})+\sin(\color{orangered}{b})\cos(\color{royalblue}{a})=\sin(\color{royalblue}{a}+\color{orangered}{b})\] to our problem:\[\Large\rm \sin(\color{royalblue}{\pi/2})\cos(\color{orangered}{\pi/7})+\sin(\color{orangered}{\pi/7})\cos(\color{royalblue}{\pi/2})=?\]

OpenStudy (anonymous):

sin((pi/2)+(pi/7))?

zepdrix (zepdrix):

Mmm ok good! Then just add them (get a common denominator).

OpenStudy (anonymous):

sin(9pi/14)?

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

Oh my gosh thank you sooooo much for this! :D